OpenStudy (anonymous):

How do you solve for X in a logirythmic equation, (0.7)^3x=2.08?

OpenStudy (anonymous):

\[3x=\frac{\ln(2.08)}{\ln(.07)}\]

OpenStudy (anonymous):

I tried changing it using the change of base rule, but it gave me the wrong answer.

OpenStudy (anonymous):

i typed is wrong, should be \[3x=\frac{\ln(2.08)}{\ln(.7)}\] \[x=\frac{\ln(2.08)}{3\ln(.7)}\]

OpenStudy (anonymous):

That does not work either, the correct answer is, -0.68. But I do not know how to figure that out without just using trial and error, I tried using those methods but it comes out wrong.

OpenStudy (anonymous):

maybe you put it in the calculator incorrectly here is the answer i wrote http://www.wolframalpha.com/input/?i=ln%282.08%29%2F%283ln%28.7%29%29

OpenStudy (anonymous):

Satellite, after your done with this..can you come back to my question?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Ok thank you very much, I just recorrected the brackets and it worked. Thanks very much.

OpenStudy (anonymous):

here